Form quadratic functions given the following information:
Zeroes at 4 and –2; f(0)=4

Answer by rothauserc(4718)

Zeroes at 4 and –2; f(0)=4
:
this means we have the product of
:
(x-4) * (x+2) = x^2 -2x -8
:
this is one of the possible quadratics, the general form would be
:
a(x^2 -2x -8)
:
we are given the point (0,4), therefore
:
a(0^2 -2(0) -8) = 4
:
-8a = 4
:
a = -1/2
:
-(1/2)(x^2 -2x -8) = (-x^2/2) +2x +4
:
*********************
f(x) = -x^2/2 +2x +4
*********************
: